OpenStudy (anonymous):

Help, will medal and fan! If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

@Loser66 will you help ?

OpenStudy (loser66):

I think if you give me some idea, I can help. ;) What is the equation of a line?

OpenStudy (anonymous):

it gives me y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3 them are the answers

OpenStudy (loser66):

the slope form of a line is y = mx +b, where m is slope. In your case, they give you m = -2, hence, it is y = -2x +b now, you need find b the line passes through (0,3), that is if x =0, y =3, the point is on the line. So, you just plug the point in y = -2x+b to find b, it is 3= -2(0) +b --> b =3 now plug back, y = -2x +3 is the required line I don't know how can you get the answer. Although it is correct but when jumping into the answer without any argument, I don't think your Prof will give you the whole credit. :)

OpenStudy (anonymous):

Thanks ! but which one is the answer then? hahha